# How to be an orthodox quantum mechanic

### v1

@@ Abstract
This work sets out to answer a single question: what is the orthodox interpretation of quantum mechanics? However, we adopt a different approach to that normally used. Rather than carefully surveying the precise details of the thoughts of Bohr and Heisenberg, we extract an orthodoxy empirically. To do this we review a collection of $33$ textbooks on quantum mechanics, encompassing the most popular and prominent works of this nature. We then gauge their response to 12 propositions to build up a picture of exactly what is believed by an orthodox quantum mechanic. We demonstrate that this orthodoxy is largely unchanged over the past century, with some interesting emerging deviations, and has many aspects of Copenhagen-like viewpoints. However, it is more nuanced than some reductive characterisations that condense it down to the ontological primacy of the quantum state. The revealed orthodoxy has two main pillars: measurement inherently disturbs quantum states and these states refer to individual instances, not ensembles. More fully it entails that individual particles exist in wave-like super-positions and present particle behaviours only when forced to by outside influences. The act of measuring such a system inherently changes its state in a random fashion, manifesting in a form of measurement error that corresponds to the uncertainty principle. This implies that measurement does not reveal underlying values of quantum properties.
@@
Quantum foundations; Quantum interpretation
@@ 1 Introduction
How do scientists understand quantum mechanics? It is commonly asserted that the most popular interpretation of quantum mechanics (QM) is the “Copenhagen interpretation”. This is, however, notoriously vague [1, 2, 3, 4] and its characterisation tends to descend into hyperfine readings of Bohr and Heisenberg [2]. Rather than characterise this Copenhagen model, it seems more productive to instead determine what the consensus within the physics community actually is. In other words, we can look for a quantum mechanical orthodoxy in the many treatises that have followed from the foundational works usually analysed to define a Copenhagen interpretation.
@@
The question of what the orthodox view of quantum theory is has, of course, received previous attention in the literature. Home [4] argues that the primary characteristic of an orthodox viewpoint is that the state is framed as a complete description of a quantum system. Thus, this encapsulates both Copenhagen-like and ensembles-based theories. This latter inclusion stems from a claim that ensemble theories are internally incomplete unless they assume completeness of the quantum state (otherwise they must supply some hidden variables to assign definite values to measurable properties). Home refers to Feyeraband’s encapsulation of the standard position as being that quantum systems possess no definite properties over and above those that can be derived from the state itself [5]. Max Jammer sets out the quantum formalism as a set of 5 postulates, all of which are necessary to an orthodox or Copenhagen-like view [3]. These include the completeness of the state and the projection postulate. To produce something like a Copenhagen interpretation this is supplemented by a kind of positivist objection to ascribing as real anything that cannot be directly measured, generating what Shimony [6] described as “doing epistemology without ontology”. This all results in a similar view to Home, that it is the primacy of the quantum state that typifies the orthodox view. Interestingly, Stapp argues [2] that a Copenhagen viewpoint does not see the quantum state as describing the evolution of the actual system in question, only its probabilities. This seems somewhat difficult to square with the notion that the state is the most complete specification of the real situation, without some explanation of why additional information is inaccessible. These points reveal a tension in descriptions of a quantum orthodoxy: a Copenhagen-like view wants to claim a “completeness” or “finality” for QM (the “end of the road hypothesis” as Popper terms it [7]). However, it also contains a strong degree of instrumentalism, as acknowledged by both Home and Jammer. This latter aspect would deny any ontological aspects to the theory. It follows then that the state is merely a prescription for the prediction of experimental outcomes. It seems clear that a follower of an allegedly Copenhagen viewpoint could pick and choose between these two, somewhat opposed, strands of thought.
Clearly, it is not sufficient to characterise the orthodoxy of quantum interpretation by the single question “is the quantum state a complete description of reality?” At least not without presupposing the logical coherency of the orthodoxy and erasing the nuance in what we mean by “complete description”. This is typified by Home’s collapse of ensemble views into the orthodoxy. We must note that there is, so far, no experiment whose statistics are not accurately predicted by quantum mechanics. However, this does not itself entail all of the views associated with Bohr and Heisenberg that both Home and Jammer link to a standard viewpoint. What we need is a more detailed breakdown of interpretative beliefs that constitute the orthodoxy, divorced carefully from the common formalism.
@@
How then do we study the beliefs of physicists as a community? The answer is that we turn to textbooks. The reason for this is that these works are presented for the education of a new generation of scientists. Thus, they are discussions of what we generally believe to be the ‘established facts’ of a given field. The key question is would the ‘established facts’ include an interpretive viewpoint? Perhaps textbooks function merely to convey a mathematical framework to their reader, with any interpretive remarks being negligible asides. This suggestion might have some validity, as a student of science can very often develop their intuition and understanding of a subject via that of the mathematics representing it. However, this is not straight-forwardly so in the realm of quantum mechanics. Evidenced by the proliferation of wildly different interpretations sharing the same (or very similar) mathematics. Thus, it is insufficient to pass on only a common mathematical core in this subject area. After all, most physics students will sketch out strikingly similar explanations of various unusual quantum results, even if their later work is done in deeply unrelated fields. They will also not merely bore a causal interlocutor with a barrage of equations. Where then would a student of quantum mechanics be exposed to an orthodoxy if not their textbooks? Perhaps in popular science media? However, one suspects that such media would itself be based on textbooks, rather than wade through the mire of competing interpretations in the active literature. An aspect of this can be seen in the fact that more recent textbooks devote more space to the interpretation of quantum phenomena (many have this as a separate chapter even). In addition, while working physicists may later deviate from the interpretations presented in reference works, we must acknowledge that their trajectory will be influenced by the initial conditions. Hence, the dominance of Copenhagen-like presentations in textbooks might lead to unorthodox interpretations which share many characteristics of this ‘reference orthodoxy’. An example of this would be the insistence that a quantum state does not apply merely to ensembles (as will be seen this is one of the most common assertions in textbooks). This is inherited by the vast majority of modern unorthodox interpretations (e.g. Bohmian mechanics [8], quantum Bayesianism [9], consistent histories [10], and relational quantum mechanics [11]), despite it not being necessary to the formalism [12]. Thus, even the beliefs of established researchers in quantum interpretation may be strongly correlated with the orthodoxy found in textbooks. It is therefore suggested that one cannot completely dismiss the interpretive remarks made in reference works on this subject. They fulfill a vital function in passing along a common understanding of quantum mechanics that is insufficiently conveyed by the mathematical formalism alone.
@@
This work attempts an empirical investigation of the quantum orthodoxy via a survey of popular and widely used textbooks on quantum mechanics, for both graduates and undergraduates. This encompasses $33$ works, with dates of first publication spanning from 1930 to 2019. The sample starts with a set of works listed as “notable” on Wikipedia [^1] https://en.wikipedia.org/wiki/List_of_textbooks_on_classical_mechanics_and_quantum_mechanics and expands it via textbooks that rank among Amazon best-sellers or are written by notable authors. To determine an orthodoxy empirically, we require some kind of framework of positions that might be part of the orthodoxy. Thus, we establish a set of 12 propositions, divided into three categories. The first of these encompasses how we understand the state in quantum mechanics. Here we identify three propositions that provide a Copenhagen-like framework: “the state is ontologically complete”, “the state does not merely describe an ensemble of identically prepared systems”, and “the state collapses upon measurement”. The second category revolves around the nature of particles with the propositions: “individual particles have wave-like properties”, “particle-like entities are described by a wave-particle duality”, “particles cannot simultaneously possess definite values of non-commuting observables”, “particles do not follow classical-like trajectories”, and “there is no adequate single picture of particle or wave”. Finally, we turn to the question of why quantum mechanics predicts statistics only. Here we have the propositions: “measurement inherently disturbs quantum systems”, “measurement creates the observed definite values”, “the uncertainty principle results from measurement disturbance”, and “the limitation to prediction of probabilities results from the uncertainty principle”. These will all be elaborated on in later sections of this paper. Each work is scrutinised and a trinary value is assigned to their agreement with each proposition (an in-between value is used when agreement can be strongly inferred from the context rather than explicitly stated). It can then be determined to what degree the works assent to each proposition on average and, by analysing the statistics, a set of propositions that robustly represent a consensus view point can be extracted (should one exist).
@@ 2 Propositions of an orthodoxy
Defining all aspects of an interpretative orthodoxy is a difficult proposition. The approach taken here is to define three sub-categories of hopefully fundamental aspects of the interpretation of the quantum formalism, these being the nature of the quantum state, that of particles, and why our predictions are all statistical. Within each category we include propositions that source works can either agree with or dissent from. The selection of these questions was designed partly apriori and partly by assertions encountered in some of the works. We can then see why there is no mention of most of the vast multitude of interpretative approaches: they tend to feature only within scientific papers, rather than in textbooks designed to educate new generations of scientists. We reiterate that the choice of textbooks over research articles is made because we feel that the most distilled consensus is to be found in textbooks, which typically are formulated around what is considered “established knowledge” within a given field. It is entirely possible that this question set is incomplete (or contains some redundancy). However, we will see that the results deduced are quite statistically robust, even with the small sample size.
@@ 2.1.1 The state is ontologically complete
This proposition indicates that there is no further information about the ‘real state of affairs’ that can be obtained once one knows the quantum state of a given system. This stands opposed to cases like the de Broglie-Bohm interpretation, where the state is supplemented by hidden variables and a pilot wave equation. An ensemble interpretation might deny this proposition, but affirm that QM is statistically complete, not because it supplements the formalism or metaphysics, rather it leaves open that QM may be a quantum analogue of statistical mechanics. Another viewpoint that would disagree with this proposition is one that assigns only epistemic value to the state. For the purposes of this work, we will take a response such as “the state represents all we can know about a system” as affirming the proposition, unless they supply some explanation of why the full ontology is not accessible.
@@ 2.1.2 The state does not just describe ensembles
Here we mean that $\psi$ describes single realisations of a given system (shorthand “single instance $\psi$”). It is not limited to the statistics of an ensemble of similarly prepared systems. This means that $\psi$ represents the current state of any single run through a given experiment. Affirming the proposition requires that individual systems can exist in indefinite/superposition states.
@@ 2.2.2 Wave-particle duality is necessary
This is an assertion of the fundamental importance of wave-like behaviour, even if the wave-like behaviour is not displayed by individual particles. Unobserved particle-like entities behave like waves, or have wave-like probability densities, but like particles when we detect them. Thus, we have dual, mutually exclusive, descriptions of the same experiments. This can be viewed as a weaker form of the preceding proposition.
@@ 2.2.5 The uncertainty principle applies to single particles
A common joke about QM goes: Heisenberg is stopped by traffic police who ask him “Do you know how fast you were going?”. His reply is “No, but I know exactly where I am”. This is the essence of this proposition. That uncertainty applies to individual realisations of quantum systems. That is, if I measure the position of a particle I cannot also know its momentum simultaneously. More generally: systems cannot posses definite values of non-commuting observables simultaneously. This is opposed to a statistical understanding which recognises that the uncertainty principle is derived via standard deviations over an ensemble [7].
@@ 2.3 Why are our predictions statistical?
When we conduct quantum mechanical experiments we require measuring over an ensemble in order to compare to the probabilistic predictions of the theory. If our interpretation of the quantum state is not ensemble-based, we have to explain why this form of measurement should be the status quo. These propositions examine some explanations (and related ideas) that arise in the studied works, in order to determine which ones may belong to the orthodoxy.
@@ 2.3.1 Measurement inherently disturbs the system
Measurement inherently changes the state of a system while measuring it (except in the case of eigenstates). A viewpoint that denies this would note that non-interactive measurements exist in QM or that measurement has no general abstract definition.
@@ 2.3.2 Uncertainty results from disturbance
The source of the uncertainty principle is the fact that our measurements must disturb quantum systems. This results in a random statistical scatter in our measurements over an ensemble of similarly-prepared systems.
@@ 2.3.3 Uncertainty leads to probabilities
This proposition regards the uncertainty principle as a fundamental feature of nature. Particularly, this also requires single-particle uncertainty. The consequences of this is that measurements have probabilistic results, even though quantum mechanics is not formulated as fundamentally statistical. One might deny this if one recognises that the uncertainty principle is a consequence of quantum rules, rather than a postulate, or if one correctly notes the uncertainty principle is formulated in terms of statistical observables only.
@@ 2.3.4 Measurement creates observed values
The observed values of quantum properties are the result of an interaction between apparatus and system, in the sense that they do not reflect underlying pre-measurement values in cases that are not eigenstates of observable being measured. This stands in opposition to the usual scientific idea of measurement: that we are generally discovering the values of pre-existing properties. This assertion can take two forms: a strong form, where it is asserted that quantum systems do not actually possess definite properties till measured, and a weaker form. In the latter, authors might take the position that measurement merely changes pre-existing values. Both of these result in the pre-measurement value being indeterminate and irrelevant.

### v2

@@ Abstract
This work sets out to answer a single question: what is the orthodox interpretation of quantum mechanics? However, we adopt a different approach to that normally used. Rather than surveying physicists, or poring over the precise details of the thoughts of Bohr and Heisenberg, we review a collection of 42 textbooks on quantum mechanics, encompassing the most popular and prominent works of this nature. We then gauge their response to 12 propositions to build up a picture of exactly what is believed by an orthodox quantum mechanic. We demonstrate that this orthodoxy has many aspects of Copenhagen-like viewpoints, but also shows some interesting emerging deviations. Moreover, it is more nuanced than some reductive characterisations that condense the orthodoxy down to the ontological primacy of the quantum state. The revealed orthodoxy has two consistent pillars: measurement inherently disturbs quantum states and these states refer to individual instances, not ensembles. More fully, it entails that individual particles exist in wave-like super-positions and present particle behaviours only when forced to by outside influences. The act of measuring such a system inherently induces random changes in its state, manifesting as a form of measurement error that corresponds to the uncertainty principle. This implies that measurement does not reveal underlying values of quantum properties.
@@
quantum mechanics, quantum interpretation, quantum foundations
@@ 1 Introduction
How do scientists understand quantum mechanics? The most popular interpretation of quantum mechanics (QM) is usually found in surveys to be “Copenhagen interpretation” [1, 2]. However, the definition of this viewpoint is notoriously vague [3, 4, 5, 6], its characterisation tending to descend into hyperfine readings of Bohr and Heisenberg [4].
@@
Max Jammer sets out the quantum formalism as a set of 5 postulates, all of which are necessary to an orthodox or Copenhagen-like view [5]. These include the completeness of the state and the projection postulate. To produce something like a Copenhagen interpretation, this is supplemented by a kind of positivist objection to ascribing as real anything that cannot be directly measured, generating what Shimony [7] described as “doing epistemology without ontology”. This all results in a similar view to Home, that it is the primacy of the quantum state that typifies the orthodox view. Interestingly, Stapp argues [4] that a Copenhagen viewpoint does not see the quantum state as describing the evolution of the actual system in question, only its probabilities. This seems somewhat difficult to square with the notion that the state is the most complete specification of the real situation, without some explanation of why additional information is inaccessible. Home [6] argues that the primary characteristic of an orthodox viewpoint is that the state is framed as a complete description of a quantum system. Thus, this encapsulates both Copenhagen-like and ensembles-based theories. This latter inclusion stems from a claim that ensemble theories are internally incomplete unless they assume completeness of the quantum state (otherwise they must supply some hidden variables to assign definite values to measurable properties). Home refers to Feyeraband’s encapsulation of the standard position as being that quantum systems possess no definite properties over and above those that can be derived from the state itself [8].
These points reveal a tension in descriptions of a quantum orthodoxy: a Copenhagen-like view wants to claim a “completeness” or “finality” for QM (the “end of the road hypothesis” as Popper terms it [9]). However, it also contains a strong degree of instrumentalism, as acknowledged by both Home and Jammer. This latter aspect would deny any ontological aspects to the theory. It follows then that the state is merely a prescription for the prediction of experimental outcomes. It seems clear that a follower of an allegedly Copenhagen viewpoint could pick and choose between these two, somewhat opposed, strands of thought.
Rather than characterise this Copenhagen model, it seems more productive to instead determine what the consensus within the physics community actually is. With a particular focus on answers to questions about quantum phenomena, along the lines of those used in [1]. Existing surveys of physicists tend to reveal a very fragmented community [1, 2], instead we will look for a quantum mechanical orthodoxy in the many treatises that have followed from the foundational works usually analysed to define a Copenhagen interpretation.
@@
The target of our survey will be textbooks. The reason for this is that these works are presented for the education of a new generation of scientists. Thus, they are discussions of what we generally believe to be the ‘established facts’ of a given field. The key question is would the ‘established facts’ include an interpretive viewpoint? Perhaps textbooks function merely to convey a mathematical framework to their reader, with any interpretive remarks being negligible asides. This suggestion might have some validity, as a student of science can very often develop their intuition and understanding of a subject via that of the mathematics representing it. However, this is not straight-forwardly so in the realm of quantum mechanics. Evidenced by the proliferation of wildly different interpretations sharing the same (or very similar) mathematics. Thus, it is insufficient to pass on only a common mathematical core in this subject area. After all, most physics students will sketch out strikingly similar explanations of various unusual quantum results, even if their later work is done in deeply unrelated fields. They will also not merely bore a causal interlocutor with a barrage of equations. Where then would a student of quantum mechanics be exposed to an orthodoxy if not their textbooks? Perhaps in popular science media? However, one suspects that such media would itself be based on textbooks, rather than wade through the mire of competing interpretations in the active literature. An aspect of this can be seen in the fact that more recent textbooks devote more space to the interpretation of quantum phenomena (many have this as a separate chapter even). In addition, while working physicists may later deviate from the interpretations presented in reference works, we must acknowledge that their trajectory will be influenced by the initial conditions. Hence, the dominance of Copenhagen-like presentations in textbooks might lead to unorthodox interpretations which share many characteristics of this ‘reference orthodoxy’. An example of this would be the insistence that a quantum state does not apply merely to ensembles (as will be seen this is one of the most common assertions in textbooks). This is inherited by the vast majority of modern unorthodox interpretations (e.g. Bohmian mechanics [10], quantum Bayesianism [11], consistent histories [12], and relational quantum mechanics [13]), despite it not being necessary to the formalism [14]. Thus, even the beliefs of established researchers in quantum interpretation may be strongly correlated with the orthodoxy found in textbooks. It is therefore suggested that one cannot dismiss the interpretive remarks made in reference works on this subject. They fulfil a vital function in passing along a common understanding of quantum mechanics that is insufficiently conveyed by the mathematical formalism alone.
This work attempts an empirical, if unavoidably subjective, investigation of the quantum orthodoxy via a survey of popular and widely used textbooks on quantum mechanics, for both graduates and undergraduates. Since we are investigating interpretation (or perhaps beliefs) this cannot be wholly objective. First, one must select which books to survey, there being no means of defining a representative sample. Our sample encompasses $42$ works, with dates of first publication spanning 1930 to 2025. The sample starts with a set of works listed as “notable” on Wikipedia [^1] https://en.wikipedia.org/wiki/List_of_textbooks_on_classical_mechanics_and_quantum_mechanics and expands it via textbooks that rank among Amazon bestsellers or are written by notable authors. This is not a rigorously chosen sample, if such a thing could even be defined in this context. However, we are searching for points of consensus, so a large sample containing the most popular works should be sufficient for our purposes. Additionally, bias from sample selection is likely to be less of an issue here than in surveys of physicists themselves.
@@
To determine an orthodoxy empirically, we require some kind of framework of positions that might be part of the orthodoxy. Once again, this injects subjectivity into our analysis, as it is not clear how we might choose these positions. We settle on a set of 12 propositions, divided into three categories. These are selected to at least cover the major assertions of a Copenhagen-like framework and show strong overlap with surveys [1]. The first category explores how we understand the quantum state, with the propositions: “the state is ontologically complete”, “the state does not merely describe an ensemble of identically prepared systems”, and “the state collapses upon measurement”. The second category revolves around the nature of particles with the propositions: “individual particles have wave-like properties”, “particle-like entities are described by a wave-particle duality”, “particles cannot simultaneously possess definite values of non-commuting observables”, “particles do not follow classical-like trajectories”, and “there is no adequate single picture of particle or wave”. Finally, we turn to the question of why quantum mechanics predicts statistics only. Here we have the propositions: “measurement inherently disturbs quantum systems”, “measurement creates the observed definite values”, “the uncertainty principle results from measurement disturbance”, and “the limitation to prediction of probabilities results from the uncertainty principle”. These will all be elaborated on in later sections of this paper. Each work is scrutinised, and a trinary value is assigned to their agreement with each proposition (an in-between value is used when agreement can be strongly inferred for consistency with other positions taken). It can then be determined to what degree the works assent to each proposition on average and, by analysing the statistics, a set of propositions that represent a consensus viewpoint can be extracted (should one exist).
@@ 2 Propositions of an orthodoxy
Defining all aspects of an interpretative orthodoxy is a difficult proposition. The approach taken here is to define three sub-categories of hopefully fundamental aspects of the interpretation of the quantum formalism, these being the nature of the quantum state, that of particles, and why our predictions are all statistical. Within each category we include propositions that source works can either agree with or dissent from. The selection of these questions was designed partly apriori and partly by assertions encountered in some works. We reiterate that the choice of textbooks over research articles is made because we feel that the most distilled consensus is to be found in textbooks, which typically are formulated around what is considered “established knowledge” within a given field. It is entirely possible that this question set is incomplete (or contains some redundancy).
We also note that our choices cover questions asked of working physicists in surveys [1, 2]. Notably, the aforementioned [1] also includes “are values are defined prior to measurement?” (“measurement creates values”), “is randomness inherent in QM?” (strong overlap with “ontological completeness”), “where is the electron in a hydrogen orbital?” (“no trajectories”), “is macroscopic superposition possible?” (linked to “the state doesn’t just describe an ensemble”), and a question on the solution of the measurement problem is covered by the entire set of questions on why QM is statistical. However, we do not include questions about the role of the observer or the implications of Bell inequality violations. These latter points are highly contentious in the literature and do not receive consistent coverage in textbooks. The cases of incomplete overlap are down to the luxury of expositive works like textbooks. In that, our questions can be simpler, more numerous, and less nuanced than those in [1, 2], as they need not be processed in real time by the respondent.
@@ 2.1.1 The state is ontologically complete
This proposition indicates that there is no further information about the ‘real state of affairs’ that can be obtained once one knows the quantum state of a given system. This stands opposed to cases like the de Broglie-Bohm interpretation, where the state is supplemented by hidden variables and a pilot wave equation. Another viewpoint that would disagree with this proposition is one that assigns only epistemic value to the state. On the other hand, an ensemble interpretation might deny this proposition, but affirm that QM is statistically complete, not because it supplements the formalism or metaphysics, rather it leaves open that QM may be a quantum analogue of statistical mechanics.
A question arises here in how we delineate epistemic viewpoints. In this work we take “the state contains all that can be known about the system” as assent to ontological completeness. The “known” aspect might make this seem questionable. However, we focus on the “can be known”, which implies an ontological limit on knowledge. If a work asserts that the state is a “calculational tool only” or that it “only characterises our knowledge of the system”, we will deem this dissent from ontological completeness.
@@ 2.1.2 The state does not just describe ensembles
Here we mean that $\psi$ describes single realisations of a given system (shorthand “non-ensemble $\psi$”). It is not limited to the statistics of an ensemble of similarly prepared systems. This means that $\psi$ represents the current state of any single run through a given experiment. Affirming the proposition requires that individual systems can exist in indefinite/superposition states.
@@ 2.2.2 Wave-particle duality is necessary
This is an assertion of the fundamental importance of wave-like behaviour, even if individual particles do not display wave behaviour. We have dual, mutually exclusive, descriptions of the same experiments. This can be viewed as a weaker form of the preceding proposition.
@@ 2.2.5 The uncertainty principle applies to single particles
A common joke about QM goes: Heisenberg is stopped by traffic police who ask him “Do you know how fast you were going?”. His reply is “No, but I know exactly where I am”. This is the essence of this proposition. That uncertainty applies to individual realisations of quantum systems. That is, if I measure the position of a particle I cannot also know its momentum simultaneously. More generally: systems cannot possess definite values of non-commuting observables simultaneously. This is opposed to a statistical understanding which recognises that the uncertainty principle is derived via standard deviations over an ensemble of measurements performed on similarly prepared systems [9].
@@ 2.3 Why are our predictions statistical?
When we conduct quantum mechanical experiments we require measuring over an ensemble of similarly prepared systems in order to compare to the probabilistic predictions of the theory. If our interpretation of the quantum state is not ensemble-based or epistemic, we have to explain why this form of measurement should be the status quo. These propositions examine some explanations (and related ideas) that arise in the studied works, in order to determine which ones may belong to the orthodoxy.
@@ 2.3.1 Measurement inherently disturbs the system
Measurement inherently changes the state of a system while measuring it (except in the case of eigenstates). A viewpoint that denies this would note that non-interactive measurements exist in QM or that measurement has no abstract definition.
@@ 2.3.2 Uncertainty results from disturbance
The source of the uncertainty principle is the fact that our measurements must disturb quantum systems. This results in a random statistical scatter in our measurements over an ensemble of similarly-prepared systems. An opposing view would be that uncertainty is present in quantum states even during Schrödinger evolution, so it cannot be associated with disturbance, or that the uncertainty principle can be validated on separate, similarly prepared samples (one set for $\hat{x}$ and the other for $\hat{p}$ for instance).
@@ 2.3.3 Uncertainty leads to probabilities
This proposition regards the uncertainty principle as a fundamental feature of nature. Particularly, this also requires single-particle uncertainty. The consequences of this is that measurements have probabilistic results, even though quantum mechanics is not formulated as fundamentally statistical. One might deny this if one recognises that the uncertainty principle is a consequence of quantum rules, rather than a postulate, or if one notes the uncertainty principle is formulated in terms of statistical observables only.
@@ 2.3.4 Measurement creates observed values
The observed values of quantum properties are the result of an interaction between apparatus and system that alters the measured property. In this sense, they do not reflect underlying pre-measurement values in cases that are not eigenstates of the observable being measured. This stands in opposition to the usual scientific idea of measurement: that we are generally discovering the values of pre-existing properties. This assertion can take two forms: a strong form, where it is asserted that quantum systems do not actually possess definite properties till measured, and a weaker form where measurement merely changes pre-existing values. Both of these result in the pre-measurement value being indeterminate and potentially irrelevant.